Grade 12Differential Calculus

An open box with a square base is to be made out of a given quantity of sheet of area a^2.
Show that: The max volume of the box is a^3/6*(3^1/2)

Faiz
12 Years agoGrade 12

1 Answer

Jitender Singh
ApprovedApproved Tutor Answer12 Years ago
Ans:
Let side of the square box is p & height is q
Volume V:
p^{2}q
Area of box:
p^{2}+pq+pq+pq+pq = p^{2}+4pq=a^{2}
q=\frac{a^{2}-p^{2}}{4p}
V=\frac{a^{2}-p^{2}}{4p}.p^{2}
V=\frac{a^{2}-p^{2}}{4}.p
\frac{\partial V}{\partial p} = 0
a^{2}-3p^{2}=0
p^{2}=\frac{a^{2}}{3}
q=\frac{a^{2}-\frac{a^{2}}{3}}{4\frac{a}{\sqrt{3}}} = \frac{a}{2\sqrt{3}}
Max. Volume:
\frac{a^{2}}{3}.\frac{a}{2\sqrt{3}} = \frac{a^{3}}{6\sqrt{3}}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty